# Euclid Quick Data Release (Q1)

### v1

@@ 5 Results for the test set
In the following, we will present the results of the analysis on the mock Euclid test set. In Sect. 5.1, we show the recovery of the parameters, along with the trends for the bias and the NMAD for each parameter. In Sect. 5.2, we analyse the behaviour of the uncertainties for each parameter. In App. A, we show the degradation in performance if one uses lenses without companions for training to predict parameters of lenses with companions in the cutouts.
@@ 5.1 Recovery of parameters for Euclid mock lenses
Figure 5: Random selection of 20 $10\arcsec\times 10\arcsec$ cutouts of simulated lenses taken from the test set. Red circles show the best predictions for the Einstein radius from LEMON, with corresponding 16th and 84th percentiles shown with dashed red circles. Yellow circles show the true values for the Einstein radii. Predictions and true values are very similar, so in some cases the corresponding circles completely overlap.
Figure 4 shows the recovery of the parameters of the Euclid mock lenses for the test set, for both the mass and light profiles, with the bottom two rows of each parameter showing the bias and NMAD metrics in sequential bins. Overall, the recovery of the parameters for the simulated Euclid lenses of the test set is good, as shown for example in Fig. 5, where we compare the Einstein radius predictions for a random selection of 20 test set lenses with the corresponding ground truth.
@@ 5.2 Behaviour of LEMON uncertainties
Figure 6: Calibrated relative uncertainty as a function of the true value of the lens parameters. For the position angle and lens magnitude, we consider the absolute uncertainty. For clarity, we plotted only $2000$ random points. The median trends of the scatter plots, along with the respective scatters associated with the 16th and 84th percentiles, are shown as red points and bars, respectively.
Figure 7: Trends of aleatoric (blue) and epistemic (orange) relative uncertainties as a function of the true value of the lens parameters. For the position angle and lens magnitude, we consider the absolute uncertainties. For clarity, we plotted only $2000$ random points from the test set.
[…]
We also verified whether the uncertainty on the parameters changes as a function of both S/N and number of companions in the test set images. For the S/N, we use the parameter $n_{\textrm{pix, source}}$, which gives the number of pixels in the lensed source image above $4\sigma$. Results show that the uncertainty for the Einstein radius slightly decreases with an increasing value of $n_{\textrm{pix, source}}$, while the uncertainties for all the other parameters slightly increase. There is, instead, no correlation between uncertainty and number of companions in the images.
The behaviour of the aleatoric and epistemic relative uncertainties associated to the test set is shown in Fig. 7, where the estimated uncertainties of both types are plotted for each lens as a function of the true values of the respective parameters. From the figure, we see that the epistemic uncertainties are systematically lower than the corresponding aleatoric uncertainties. This is the expected behaviour: if the training set correctly spans all the parameter space, the epistemic uncertainties should be as close to zero as possible.
@@ 6.1 LEMON applied to Euclidised lenses
We now apply LEMON to a sub-sample of real Euclidised lenses described in Sect. 2.2. This sub-sample is formed by four different main catalogues, from which the ground truth values have been taken (where available).
  • A sample of 29 Sloan Lens Advanced Camera for Surveys (SLACS) catalogue systems, described in Bolton et al. (2008) and Auger et al. (2009).
  • The sample of 13 early-type/early-type lens systems (EELs) presented in Oldham et al. (2017).
  • A sample of five lenses found in the COSMOS survey (Scoville et al. 2007), detailed in Faure et al. (2008).
@@
  • A sample of 13 lenses taken from those found by visually inspecting the whole imaging data taken with the ACS through the F814W filter up to the 31st of August 2011 (Pawase et al. 2014). Notice that the Einstein radius for these lenses is not reported, and as such we used the radius of the arc (in arcseconds) as a substitute.
@@
All of the Euclidised lenses used in this work have been spectroscopically confirmed. The recovery plot for these lenses is shown in Fig. 8, while the cumulative metrics for each parameter are reported in Table 3. A panel with some of the lenses from each catalogue is shown in Fig. 9.
Figure 8: Recovery plot for the parameters of the Euclidised lenses, showing how well the predicted model parameters from LEMON ($y$ coordinate) reproduce the corresponding true value ($x$ coordinate), for each lens. The ideal recovery line is shown as a dashed black line. The SLACS and ACS best values for the Sérsic indexes are evaluated as the median of the predictions against the value for a De Vaucouleurs profile ($n_{\textrm{lens}}=4$), with uncertainty given by the NMAD of the predictions. Lenses in panel (d) with true values bigger than $2\arcsec$ have been removed due to being outside the training space of LEMON.
@@
Figure 9: Random selection of 20 $10\arcsec\times 10\arcsec$ cutouts of Euclidised lenses taken from the SLACS, EEL, COSMOS, and ACS subsamples. Red circles show the best predictions for the Einstein radius from LEMON, with corresponding 16th and 84th percentiles shown with dashed red circles. Yellow circles show the values for the Einstein radii reported in Bolton et al. (2008), Oldham et al. (2017), Faure et al. (2008), and Pawase et al. (2014), respectively. For ACS lenses, the radius of the arc is taken instead, due to lack of Einstein radius measurement.
Figure 10: Distributions of the Einstein radius (left panel) and the lens effective radius (right panel) as a function of lens redshift. Black regions show the training set, while coloured points show the position of the Euclidised lenses.
@@
Table 3: Cumulative metrics associated to the recovery of the parameters of the Euclidised lenses.
[…]
Bias ($\mu$) 0.17 (0.14) 0.02 (0.01) --1.93 (--2.36) 0.19 (0.22) 0.47 (0.96)
Lower discr. scatter ($\sigma^{-}$) 0.19 (0.12) 0.10 (0.08) 11.18 (11.01) 0.51 (0.53) 1.30 (1.48)
Upper discr. scatter ($\sigma^{+}$) 0.39 (0.24) 0.10 (0.12) 12.77 (9.60)0 0.13 (0.11) 2.57 (2.52)
RMSE 0.63 (0.48) 0.11 (0.11) 32.47 (26.51) 0.38 (0.39) 1.94 (2.13)
MAE 0.42 (0.30) 0.09 (0.08) 18.10 (13.55) 0.32 (0.32) 1.46 (1.63)
NMAD 0.23 (0.14) 0.11 (0.11) 09.11 (10.13) 0.23 (0.22) 1.54 (2.06)
Coefficient of determination ($R^{2}$) --0.03 (0.91)0 0.51 (0.76) 0.57 (0.98) 0.49 (0.53) --0.16 (--0.02)
 [^7] Values in parentheses correspond to the values of the cumulative metrics on the subset of lenses with predicted low uncertainty on the Einstein radius ($\sigma_{R_{\textrm{Ein}}}\leq 0.5\arcsec$).
[…]
In order to try and reduce the noise induced on the cumulative metrics by lenses with low S/N, we re-evaluated them by removing all lenses having predicted uncertainty on the Einstein radius bigger than $0.5″$. Results are shown in the parentheses of Table 3. We can see that the metrics of Einstein radius, axis ratio, and position angle improve substantially, with the values of $R^{2}$ reaching values of up to $0.98$ for the latter. There is, however, no improvement on the metrics of the other parameters. When we remove uncertain effective radii or Sérsic indices, though, the statistical indicators for these parameters do not improve significantly.
[…]
We also checked the distribution of the Einstein and effective radii as a function of the lens redshifts, as can be seen in Fig. 10. From these, we can see that the parameters for the Euclidised lenses are often on the edge of the training-set distribution, if not completely outside it. This could influence negatively the recovery of these parameters, given that the network is not trained to recognise this region of the parameter space. In particular, in panel (d) of Fig. 8, we removed lenses with true effective radii larger than $2\arcsec$, since they fall outside the training sample of LEMON, thus making their predicted values unreliable extrapolations. Furthermore, the true estimates of the effective radii cannot always be compared homogeneously with the predictions, as SLACS and ACS lenses are modelled with a fixed De Vaucouleurs profile (i.e., $n_{\textrm{lens}}=4$), while LEMON considers Sérsic profiles with variable values of $n_{\mathrm{lens}}$. Additionally, the fitting process should ideally remove the arc to obtain an accurate estimate; otherwise, the Sérsic fit may overestimate the radius due to bumps in the surface brightness profile. In future works, we will expand the distribution of the training set to cover more of the parameter space, and check if the recovery of these parameters improves.
@@
Figure 11: $10\arcsec\times 10\arcsec$ cutouts of the five Euclid gravitational lenses found in the Perseus ERO field and modelled in Acevedo Barroso et al. (2024). Top row: unedited cutouts centred on the lenses. Bottom row: the same cutouts, with the predicted Einstein radius (red circles, with dashed circles showing the uncertainty bands) and the value obtained from the classical modelling (yellow circle) superimposed on them.
@@ 6.2 LEMON applied to real Euclid ERO lenses
We also applied LEMON to real Euclid lenses found in the Perseus ERO field and modelled in Acevedo Barroso et al. (2024). The modelling has been performed with the pronto software (Vegetti & Koopmans 2009; Rybak et al. 2015a, b; Rizzo et al. 2018; Ritondale et al. 2019; Powell et al. 2021), with the lens mass distribution modelled by an SIE and the light distribution modelled as a composite of three Sérsic profiles. The surface brightness distribution of the source is reconstructed using a pixelated, free-form approach. Regularisation is applied to penalise large gradients in the reconstructed source brightness. The parameters for the lens mass and light profiles are optimised nonlinearly using the MultiNest algorithm (Feroz et al. 2009). The positions of assumed lens images are provided as inputs, with models constrained to ensure that these positions align in the source plane with a tolerance of 11″1  ″. Each model is checked against three criteria: the presence of an SIE critical curve consistent with observed lens image; the alignment of the critical curve with the lens galaxy’s light profile; and finally, the consistency of the reconstructed source with a compact object inside a caustic.
From all the lenses found in the field, we considered the ones that have a convincing model (identified as ‘valid’ in the ‘modelling’ column of table 2 in Acevedo Barroso et al. 2024), and for which the value of the Einstein radius is available. An image of these lenses is shown in Fig. 11. The recovery plot associated to these lenses is shown in Fig. 12. From the figure, we can see that the recovery of the Einstein radius is within $1\sigma$ from the ideal equality line for all lenses, except for the lens J031959+414229. Overall, the recovery of the parameter is sufficiently accurate for real Euclid lenses. It is remarkable that the lens J031749+420011 is recovered correctly, albeit with a large uncertainty, due to the fact that there is a very bright star present in the cutout, which increases both the aleatoric uncertainty and the epistemic uncertainty, the former due to low visibility of the lens and the latter due to not having trained LEMON with stars as contaminants in the image.
Figure 12: Same as Fig. 8, but for the Euclid lenses found in the Perseus ERO field.

### v2

@@ 5 Results for the test set
In the following, we present the results of the analysis on the mock Euclid test set. In Sect. 5.1, we show the recovery of the parameters, along with the trends for the bias and the NMAD for each parameter. In Sect. 5.2, we analyse the behaviour of the uncertainties for each parameter.
@@ 5.1 Recovery of parameters for Euclid mock lenses
Figure 5: Random selection of 20 $10\arcsec\times 10\arcsec$ cut-outs of simulated lenses taken from the test set. Red circles show the best predictions for the Einstein radius from LEMON, with corresponding 16th and 84th percentiles shown with dashed red circles. Yellow circles show the true values for the Einstein radii. Predictions and true values are very similar, so in some cases the corresponding circles completely overlap.
Figure 6: Calibrated relative uncertainty as a function of the true value of the lens parameters. For the ellipticity components, lens magnitude, and source co-ordinates, we consider the absolute uncertainty. For clarity, we plotted only $5000$ random points. The median trends of the scatter plots, along with the respective scatters associated with the 16th and 84th percentiles, are shown as red points and bars, respectively.
Figure 4 shows the recovery of the parameters of the Euclid mock lenses for the test set, for both the mass and light profiles, with the bottom two rows of each parameter showing the bias and NMAD metrics in sequential bins. Overall, the recovery of the parameters for the simulated Euclid lenses of the test set is good, as is shown for example in Fig. 5, where we compare the Einstein radius predictions for a random selection of 20 test set lenses with the corresponding ground truth.
The Einstein radius is well recovered, with a very slight increase in NMAD and decrease in bias from 1.0 upwards, due to a skewed prior towards zero for the Einstein radius parameter, as we can see from Fig. 4a. The mass ellipticity components are well recovered, showing a slight increase in bias only for values outside the interval [0.5,0.5][-0.5,0.5] and a NMAD trend that always stays below $0.1$. The recovery is even tighter for the light ellipticity, showing an $R^{2}$ of $0.96$ (compared against an $R^{2}$ of $0.85\textrm{--}0.86$ for the mass ellipticity components) and a NMAD trend that always lies below $0.04$. This slight discrepancy is to be expected, because the network cannot rely on light information to infer the shape of the mass distribution, and has instead to rely only on the configuration of the lensed source to derive the mass shape parameters, which is an intrinsically more uncertain procedure.
@@ 5.2 Behaviour of LEMON uncertainties
Figure 6 shows the trends for the calibrated total relative uncertainties [^8] We define the total relative uncertainty as the predicted total uncertainty by LEMON divided by the corresponding predicted mean value. for all predicted parameters (except for the ellipticity components and lens $I_{\scriptscriptstyle\rm E}$ magnitude, where we consider the absolute uncertainty). From the figure, we can see that the relative uncertainty on the Einstein radius is of order $\sim 5\times 10^{-2}$ on all the relevant range of the parameter, with a constant scatter throughout the interval range. Regarding the mass ellipticity components, we can see that the uncertainty increases for $\epsilon_{\textrm{m,l}}\rightarrow\pm 0.8$, probably due to the fact that the distribution of the ellipticities is peaked around zero and thus there are fewer examples on the edges of the parameters. A similar trend is visible for the light ellipticity components, but with much smaller uncertainties. This reflects the fact that the network has an easier time predicting the shape of the light component directly from the lens galaxy image, compared to inferring the shape of the mass component through the lensed images’ configurations. The trend for the lens effective radius is similar to the one for the Einstein radius, with an increase at low values due to the predicted values of the parameter approaching zero. The scatter of the Srsic index is systematically high for all values of $n_{\textrm{lens}}$, probably due to an intrinsic difficulty in recovering this parameter through photometry alone. The relative uncertainty for this parameter, however, mainly remains below $40\%$. For both low and high values of lens magnitude, we find that the absolute uncertainty slightly increases. Finally, we note that the uncertainties in the unlensed source co-ordinates exhibit trends similar to those of the ellipticity components.
We also compared the trends of aleatoric and epistemic uncertainties for the same parameters of Fig. 6. We verified that the epistemic uncertainties are systematically lower than the corresponding aleatoric uncertainties for all parameters. The comparison, along with a brief discussion, is shown in Fig. 20 of Appendix D. Additionally, we analysed the behaviour of total uncertainty with the S/N of the source in Appendix E, which gave us information about how each parameter is recovered by the network (through either the light or the lensed source configuration).
@@ 6 Results for real lenses
Figure 7: Predicted model parameters from LEMON ($y$ co-ordinate) against the corresponding model values from the literature ($x$ co-ordinate). The ideal recovery line is shown as a dashed black line. The SLACS and ACS best predicted values for the Srsic indices are evaluated as the median of the predictions for all the lenses of these catalogues against a fixed value associated with the one for a De Vaucouleurs profile ($n_{\textrm{lens}}=4$), with uncertainty given by the NMAD of the predictions. Lenses with modelling values for the effective radius bigger than $2\arcsec$ and EEL lenses with modelling values for the Einstein radius lower than 0.5 have been removed due to being outside the training space of LEMON.
@@ 6.1 LEMON applied to Euclidised lenses
We now apply LEMON to a sub-sample of real Euclidised lenses described in Sect. 2.2. This sub-sample is formed from four different main catalogues, from which the reference parameter values from standard modelling have been taken (where available).
  • A sample of 29 Sloan Lens Advanced Camera for Surveys (SLACS) catalogue systems, described in Bolton et al. (2008) and Auger et al. (2009). Einstein radius, axis ratio, and position angle (estimated by assuming a SIE mass model for the lenses without external shear, and then used to derive the corresponding mass ellipticity components) are taken from the first three columns of table 5 in Bolton et al. (2008), after correcting the position angle orientation from east of north to north of west; the effective radius and lens magnitude are instead taken from the columns named $r_{{\rm e},I}$ and $m_{I}$ of table 3 in Auger et al. (2009).
  • The sample of 13 early-type source with early-type lens systems (EELs) presented in Oldham et al. (2017). Einstein radius, mass axis ratio, and mass position angle values are all taken from table 2, after the appropriate orientation conversions for the position angle. The mass model considered by the authors is an elliptical power law with external shear.
  • A sample of five lenses found in the COSMOS survey (Scoville et al. 2007). All parameters are taken from table 2, except for the Einstein radius, which is taken from the errata by Faure et al. (2008) due to an error in the determination of the mass parameters in Scoville et al. (2007). To model the systems, the authors use the code Lenstool, assuming a SIE model with external shear. Light ellipticity components have been derived from the parameters 1q1-q and PA of table 2, and used for comparison.
@@
  • A sample of 13 lenses taken from those found by visually inspecting the whole imaging data taken with the ACS through the F814W filter up to the 31st of August 2011 (Pawase et al. 2014). All parameters are taken from table 3, and light axis ratio and position angle have been converted to light ellipticity components for comparison. Notice, however, that the Einstein radius for these lenses is not reported, and as such we compared with the radius of the arc (in arcseconds) as a substitute.
@@
All of the Euclidised lenses used in this work have been spectroscopically confirmed. The comparison between forward modelling and LEMON predictions for mass and light parameters is shown [^9] Notice that we aggregated mass and light ellipticity comparisons together. Mass and light ellipticity results from the literature have been compared with the matching type of parameter prediction from LEMON. in Fig. 7, while the associated metrics for each parameter are reported in Table 3. A panel with some of the lenses from each catalogue is shown in Fig. 8.
@@
Figure 8: Random selection of 20 $10\arcsec\times 10\arcsec$ cut-outs of Euclidised lenses taken from the SLACS, EEL, COSMOS, and ACS subsamples. Red circles show the best predictions for the Einstein radius from LEMON, with corresponding 16th and 84th percentiles shown with dashed red circles. Yellow circles show the values for the Einstein radii reported in Bolton et al. (2008), Oldham et al. (2017), Faure et al. (2008), and Pawase et al. (2014), respectively. For ACS lenses, the radius of the arc is taken instead, due to lack of an Einstein radius measurement.
@@
Table 3: Cumulative metrics associated with the parameters of the Euclidised lenses shown in Fig. 7.
Metrics $R_{\textrm{Ein}}$ (arcsec) $\epsilon_{x}$ $\epsilon_{y}$ $R_{\textrm{e, lens}}$ (arcsec) $n_{\textrm{lens}}$ $m_{\textrm{lens}}$
Bias ($\mu$) $3\hphantom{\;\;\,} 0.00 0$.00 0.03 $35\hphantom{\;\;\,} 0.0$1
RMSE 0.14 0.07 0.04 0.18 1.05 0.30
NMAD 0.11 0.04 0.03 0.06 1.10 0.21
Coefficient of determination ($R^{2}$) 0.53 0.87 0.88 0.64 $0.47\hphantom{\;\;}$ 0.69
Figure 9: Distributions of the Einstein radius (left panel) and the lens effective radius (right panel) as a function of lens redshift. Black regions show the training set, while coloured points show the position of the Euclidised lenses. Notice that these points are predictions from forward modelling.
From the table, we can see that the Einstein radius predictions match well the literature results, with a coefficient of determination of $R^{2}=0.53$. The mass and light ellipticities also match the values from the literature well, with $R^{2}=0.87\textrm{--}0.88$. The effective radius is slightly underpredicted with respect to classical modelling above 1. This, however, is consistent with results from Li et al. (2022), in which light parameters of galaxies were predicted with CNNs and compared with classical modelling predictions from 2DPHOT (La Barbera et al. 2008). In both cases, there is a systematic underprediction at high values of effective radii. Due to a low number of points available for the Srsic index parameter, in addition to the fact that for SLACS and ACS lenses the classical modelling has been performed by assuming a fixed $n_{\textrm{lens}}=4$ value for this parameter, we cannot draw any particularly relevant conclusions. Finally, the lens magnitude for the SLACS lenses matches well the results from classical modelling, with $R^{2}=0.69$.
@@
Figure 10: $10\arcsec\times 10\arcsec$ cut-outs of the five Euclid gravitational lenses found in the Perseus ERO field and modelled in Acevedo Barroso et al. (2025). Top row: Unedited cut-outs centred on the lenses. Bottom row: Same cut-outs, with the predicted Einstein radius (red circles, with dashed circles showing the uncertainty bands) and the value obtained from the classical modelling (yellow circle) superimposed on them.
Figure 11: Comparison between Einstein radii values predicted by LEMON and those predicted by classical modelling for the Euclid lenses found in the Perseus ERO field.
To analyse the behaviour of LEMON with respect to the Euclidised lenses, we first evaluated both the aleatoric and epistemic components of the uncertainties for these lenses, obtaining values of epistemic uncertainties much lower than the aleatoric ones (e.g. the maximum epistemic uncertainty for the Einstein radius is at most 0.25, while the corresponding maximum aleatoric uncertainty is 1.35), in agreement with the results for the mock Euclid lenses’ test set. These values, however, are much higher than the corresponding components of the uncertainties for the mock Euclid lenses’ test set (e.g. the maximum aleatoric uncertainty for the Einstein radius for the mock Euclid lenses’ test set is 0.20, which is of the order of the maximum epistemic uncertainty for the Euclidised lenses set).
Figure 12: Parameter estimates of Euclid lenses found in the Q1 fields given by LEMON against values obtained by traditional modelling with PyAutoLens. For each panel, we also report the respective cumulative metrics. Notice that predicted LEMON magnitudes have been subtracted by a factor of $0.22$ due to different zero-points between mock Euclid images used for training and real Q1 images.
A possible cause of these high uncertainty values is distributional shift, i.e. variations in the properties of the input image shift the parameters’ distributions outside the training distribution space, which affects the network’s output predictions (for more information, see Filipp et al. 2025). To check this, we analysed the distribution of the Einstein and effective radii as a function of the lens redshifts, as can be seen in Fig. 9. From these, we can see that the parameters for the Euclidised lenses are often on the edge of the training-set distribution, if not completely outside it. This could influence negatively the recovery of these parameters, given that the network is not trained to recognise this region of the parameter space. To try and alleviate this issue, we preventively removed from our comparisons lenses with values of the Einstein radius and effective radius estimated from the literature smaller than 0.5 and larger than $2\arcsec$, respectively, since they fall outside the training sample of LEMON, making their predicted values unreliable extrapolations. Another performance driver could be the fact that the estimates of the effective radii from classical modelling cannot always be compared homogeneously with LEMON’s predictions, as SLACS and ACS lenses are modelled with a fixed De Vaucouleurs profile (i.e. $n_{\textrm{lens}}=4$), while LEMON considers Srsic profiles with variable values of $n_{\mathrm{lens}}$. Additionally, the fitting process should ideally remove the arc to obtain an accurate estimate; otherwise, the Srsic fit may overestimate the radius due to bumps in the surface brightness profile. A final possible performance driver could be the fact that the Euclidisation procedure degrades the original image, and as such faint arcs or counterimages could become invisible after the procedure. This could produce lenses that would not be detectable with Euclid resolution, and that would thus have lensing configurations that were not visible or that were outside the training space of LEMON. In future works, we shall expand the distribution of the training set to cover more of the parameter space, and check if the recovery of these parameters improves.
@@ 6.2 LEMON applied to real Euclid ERO lenses
We also applied LEMON to real Euclid lenses found in the Perseus ERO field and modelled in Acevedo Barroso et al. (2025). The modelling was performed with the pronto software (Vegetti & Koopmans 2009; Rybak et al. 2015a, b; Rizzo et al. 2018; Ritondale et al. 2019; Powell et al. 2021), with the lens mass distribution modelled by a SIE and the light distribution modelled as a composite of three Srsic profiles. The surface brightness distribution of the source was reconstructed using a pixelated, free-form approach. Regularisation was applied to penalise large gradients in the reconstructed source brightness. The parameters for the lens mass and light profiles were optimised non-linearly using the MultiNest algorithm (Feroz et al. 2009). The positions of assumed lens images were provided as inputs, with models constrained to ensure that these positions align in the source plane with a tolerance of 1. Each model was checked against three criteria: the presence of a SIE critical curve consistent with observed lens image; the alignment of the critical curve with the lens galaxy’s light profile; and finally, the consistency of the reconstructed source with a compact object inside a caustic.
From all the lenses found in the field, we considered the ones that have a convincing model (identified as ‘valid’ in the ‘modelling’ column of table 2 in Acevedo Barroso et al. 2025), and for which the value of the Einstein radius is available. An image of these lenses is shown in Fig. 10. The comparison between LEMON predictions for the Einstein radius and those from classical modelling associated with these lenses is shown in Fig. 11. From the figure, we can see that the recovery of the Einstein radius is within $1\sigma$ from the ideal equality line with respect to traditional modelling results for all lenses, except for the lens J031741+412702, for which the Einstein radius is slightly underestimated. It is remarkable that the lens J031749+420011 is recovered correctly, albeit with a large uncertainty, due to the fact that there is a very bright star present in the cut-out, which increases both the aleatoric uncertainty and the epistemic uncertainty, the former due to low visibility of the lens and the latter due to not having trained LEMON with stars as contaminants in the image.